OpenStudy (anonymous):

HELP!! the area of a regular polygon is 480 and its apothem is 20. find the perimeter

OpenStudy (e.mccormick):

The area of any regular polygon is: \[\frac{1}{2}\times apothem\times perimeter\]You have the area and apothem, so you can turn this into:\[480=\frac{1}{2}(20)P\] Now solve for p.